I think I solved this correctly, but just wanted to check:
9y+3x=18; -3y-x=-6
9y=18-3x
y=2-1/3x
Substitution for -3y-x=-6
-3(2-1/3x)-x=-6
-6-x-x=-6
-6-2x=-6
x=6
9y+3(6)=18
9y=0
y=0
We are supposed to solve for either variable and use the solution to determine the second variable.
I got x=6, y=0, but am unsure if this is correct.

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

9y+3x=18 --> x + 3y = 6
-3y-x=-6 --> x + 3y = 6
---------------
It's the same eqn twice
--> no unique solution, an infinite number of solutions.

Answer by rothauserc(4718)

you are correct!
best way to check your answers is to substitute values for x and y in the original equations
x = 6, y = 0
9*0 +3*6 = 18
18 = 18
answers in equation 1
-3y -x = -6
-3*0 -6 = -6
-6 = -6
answers check in equation 2